OpenStudy (anonymous):

unsimplified solution for x2 + 8x + 6 = 0? WILL MEDAL AND FAN

OpenStudy (anonymous):

no, it's in a form like \[x=-8+\sqrt{40}\\] -------------------- 4

jimthompson5910 (jim_thompson5910):

Use the Quadratic Formula \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

in this case a = 1 b =8 c = 6

OpenStudy (anonymous):

yeah the answer's -7.16, -.83 i think but idk how to show it unsimplified.

jimthompson5910 (jim_thompson5910):

well instead of evaluating, you just stop beforehand

jimthompson5910 (jim_thompson5910):

what is the value of \[\Large b^2-4ac\]

OpenStudy (anonymous):

40

jimthompson5910 (jim_thompson5910):

b = 8, so -b = -8 a = 1, so 2a = 2*1 = 2

jimthompson5910 (jim_thompson5910):

that means we go from \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] to \[\Large x=\frac{-8 \pm \sqrt{40}}{2}\]

OpenStudy (anonymous):

ohhhh okay thanks

jimthompson5910 (jim_thompson5910):

np